### v4

# Inverse problems related to electrical networks and the geometry of non-negative Grassmannians

Abstract.
We provide a new solution to the classical black box problem (the discrete Calderón problem) in the theory of circular electrical networks. Our approach is based on the explicit embedding of electrical networks into non-negative Grassmannians and generalized chamber ansatz for it. Also, we reveal the relation of this problem with the combinatorial properties of spanning groves and the theory of totally non-negative matrices.
MSC2020: 05C10, 05C50, 05E10, 15M15, 34B45, 35R30, 82B20, 90C05, 90C59, 94C05, 94C15
Key words: electrical networks, discrete Calderón problem, discrete electrical impedance tomography, non-negative Grassmannians, twist for positroid variety, Temperley trick, totally non-negative matrices, effective resistances.
1. Introduction
1.1. Organization of the paper
An electrical network $\mathcal{E}(G,w)$ consists of a planar graph $G$ equipped with positive edge weights $w$ denoting their conductivities. The nodes of $\mathcal{E}(G,w)$ are partitioned into two sets: inner nodes and boundary nodes. By applying voltages $\textbf{U}:V_{B}\to\mathbb{R}$ at the boundary nodes and using Ohm’s and Kirchhoff’s laws, the voltages at all nodes $U:V\to\mathbb{R}$ are uniquely determined. These voltage extensions give rise to two natural network characteristics: the electrical response matrix and the effective resistance matrix.
Combining the generalized Temperley trick with the relation between the entries of these matrices and spanning forests enables the construction of an embedding into the Grassmannian $Gr_{\geq 0}(n-1,2n)$, see Sections 2.2 and 3.2. This embedding can be explicitly defined via response matrices or effective resistance matrices, see Theorem 2.14 and Theorem 2.14. That allows us to apply the Postnikov theory (see Section 2.1) of non-negative Grassmannians to the study of electrical networks. Particularly, drawing upon the Postnikov theory, we propose (see Section 3) a novel algorithmic solution to the black box problem, which is the main result of our paper (note that a similar approach has been independently studied in [23]):
Proposition 1.1.
Consider a minimal electrical network $\mathcal{E}(G,w)$ and its associated point $\mathcal{L}(\mathcal{E})$ of non-negative Grassmannian $Gr_{\geq 0}(n-1,2n)$. Then, each edge conductivity can be expressed as monomial ratios of a special set of Plücker coordinates of the twisted point $\tau\bigl(\mathcal{L}(\mathcal{E})\bigr)$ of $Gr_{\geq 0}(n-1,2n)$ corresponding to weights of minimal spanning groves. The indices of these Plücker coordinates are determined by applying the Scott rule to the minimal Lam model, which is constructed from $\mathcal{E}(G,w)$ using the generalized Temperley trick.
Unlike other existing methods, our solution is applicable for any network and has a non-recursive nature, which we anticipate might lead to more numerically stable approaches to the Calderón problem (see Sections 1.3) and some problems in phylogenetic network theory, see references in Section 3.2.
Finally, in Section 4 we reveal the principal relation of the black box problem to the classical theory of totally non-negative matrices.
Acknowledgments. The work on Sections 1, 2, 3 was funded by the Russian Science Foundation project No. 20-71-10110 (https://rscf.ru/en/project/23-71-50012) which finances the work of the author at P. G. Demidov Yaroslavl State University. The work on Sections 4 was supported by the Ministry of Science and Higher Education of the Russian Federation (Agreement No. 075-02-2025-1636).
The author is grateful to Vassily Gorbounov, Dmitry Talalaev, Sotiris Konstantinou-Rizos, and Lazar Guterman for their invaluable discussions and careful reading of this work. Special thanks are due to Lazar Guterman for his assistance with the figures.
1.2. Electrical network theory
We start with a brief introduction into the theory of electrical networks following [7], [8] and [14].
Definition 1.2.
An electrical network $\mathcal{E}(G,w)$ is essentially a planar graph $G(V,E)$, embedded into a disk and equipped with a conductivity function $w:E(G)\to\mathbb{R}_{>0}$, which satisfies the following conditions:
  • All nodes are divided into the set of inner nodes $V_{I}$ and the set of boundary nodes $V_{B}$ and each boundary node lies on the boundary circle;
  • Boundary nodes are enumerated clockwise from $1$ to $|V_{B}|:=n.$ Inner nodes are enumerated arbitrarily from $n+1$ to $|V|$;
  • An edge weight $w(v_{i}v_{j})=w_{ij}$ denotes the conductivity of this edge.
Consider an electrical network $\mathcal{E}(G,w)$ and apply voltages $\textbf{U}:V_{B}\to\mathbb{R}$ to its boundary nodes $V_{B}$. Then these boundary voltages induce the unique harmonic extension on all vertices $U:V\to\mathbb{R},$ which might be found out by Ohm’s and Kirchhoff’s laws:
$\sum\limits_{j\in V}w_{ij}\bigr(U(i)-U(j)\bigl)=0,\ \forall i\in V_{I}.$
One of the main objects associated with each harmonic extension of boundary voltages U are boundary currents $\textbf{I}=\{I_{1},\dots,I_{n}\}$ running through boundary nodes:
$I_{k}:=\sum\limits_{j\in V}w_{ij}\bigr(U(k)-U(j)\bigl),\ k\in\{1,\dots,n\}.$
The first cornerstone result in the theory of electrical networks is that boundary voltages and currents are related to each other linearly.
Theorem 1.3.
[8] Consider an electrical network $\mathcal{E}(G,w).$ Then, there is a matrix $M_{R}(\mathcal{E})=(x_{ij})\in Mat_{n\times n}(\mathbb{R})$ such that the following holds:
$M_{R}(\mathcal{E})\textbf{U}=\textbf{I}.$
This matrix is called the response matrix of a network $\mathcal{E}(G,\omega)$ and it satisfies the following conditions:
  • Matrix $M_{R}(\mathcal{E})$ is symmetric;
  • All non-diagonal entries $x_{ij}$ of $M_{R}(\mathcal{E})$ are non-positive;
  • For each row ((column)) the sum of all its entries is equal to $0.$
  • Matrix $M_{R}(\mathcal{E})$ is circular totally non-negative i.e. for each $k\in\{1,\dots,n\}$ all $k\times k$ circular minors of $M_{R}(\mathcal{E})$ ((see [8] and [14] for more details)) are signed non-negative: $(-1)^{k}\det M_{R}(\mathcal{E})(P;Q)\geq 0,$ for each circular pair $(P;Q).$
The following five local network transformations given below are called the electrical transformations.
Figure 1. Electrical transformations
Two electrical networks are defined to be equivalent if they can be obtained from each other by a sequence of electrical transformations. The importance of these transformations is revealed by the following result.
Theorem 1.4.
[7] The electrical transformations preserve the response matrix of an electrical network. Any two electrical networks that share the same response matrix are equivalent.
According to Theorem 1.4, we will consider electrical networks up to electrical transformations. We will denote by $E_{n}$ the set of all electrical networks with $n$ boundary nodes up to electrical transformations. In fact, each electrical network can be transformed into a minimal network.
Definition 1.5.
A median graph of an electrical network $\mathcal{E}(G,w)$ is a graph $G_{M}$ whose internal vertices are the midpoints of the edges of $G$ and two internal vertices are connected by an edge if the edges of the original graph $G$ are adjacent. The boundary vertices of $G_{M}$ are defined as the intersection of the natural extensions of the edges of $G_{M}$ with the boundary circle. Since the interior vertices of the median graph have degree four, we can define the strands of the median graph as the paths which always go straight through any degree four vertex.
Figure 2. Star-shape network and its median graph
Definition 1.6.
An electrical network is called minimal if the strands of its median graph do not have self-intersections; each strand does not form a closed loop; any two strands intersect at most once i.e. the median graph has no lenses, see Fig. 3.
Figure 3. A lens obtained by the intersection of strands $\alpha_{1}$ and $\alpha_{2}$
Theorem 1.7.
[8],[14]
  • Each electrical network $\mathcal{E}(G,w)\in E_{n}$ is equivalent to a minimal network;
  • Any two minimal electrical networks which share the same response matrix can be transformed to each other only by the star-triangle transformation.
1.3. The black box problem for electrical networks
One of the most important and notoriously difficult problems in electrical network theory, which has significantly influenced the entire development of this field [8], is known as the black box problem.
Problem 1.8.
Consider an electrical network $\mathcal{E}(G,w)$ on a given graph $G.$ The conductivity function $w$ is required to be recovered by a known response matrix $M_{R}(\mathcal{E}).$
The following theorem reveals the possibility of solving the black box problem:
Theorem 1.9.
[7],[8] The conductivity function $w$ of a minimal electrical network $\mathcal{E}(G,w)\in E_{n}$ can be uniquely recovered by its response matrix $M_{R}(\mathcal{E}).$
Notably, all proofs of Theorem 1.9 are constructive and directly provide practical algorithms for solving Problem 1.8. We will now highlight a few of the main ones:
  • The recursive spike-bridge-pendant decomposition [8];
  • Solutions to Problem 1.8 for special (for instance, standard) electrical networks [8], [13];
  • Reducing Problem 1.8 to the black box problem in the Postnikov theory. This approach will be described in Section 2 (also see [23]).
Problem 1.8 plays a central role in electrical network theory because it enables the development of fast and stable numerical solutions to the well-known Calderón problem [24], [25]. In fact, Problem 1.8 can be viewed as a discrete analog [3] of a specific instance of the Calderón problem, known as the inverse problem of electrical impedance tomography in a circle. For completeness, we will briefly outline the formulation of this problem and its link to Problem 1.8.
Consider a conducting medium in a closed disk $D$ i.e. a positive bounded function of conductivity $\sigma(x):D\to\mathbb{R}_{>0}$ and an electric voltage $u(x):D\to\mathbb{R}$, which are related to each other by the following second-order elliptic partial differential equation:
$div(\sigma(x)\nabla u(x))=0,\ \forall x\in int\ D,$
with Dirichlet boundary data which is a known voltage $V:\partial D\to\mathbb{R}$ applied to the boundary of $D$:
$u(x)_{|\partial D}=V(x),\ \forall x\in\partial D.$
The voltage $V(x)$ induces a Neumann boundary data, namely currents $I_{V}(x)$ running through $\partial D$$\partial D\to\mathbb{R}$ induced by $V(x)$ :
$I_{V}(x)=(n,\sigma(x)\nabla u(x)_{|\partial D}),$
where $n$ is the unit vector of the outer normal to $\partial D;$ and $(\cdot,\cdot)$ is the standard scalar product.
Suppose that for each boundary voltage $V(x)$ the corresponding induced boundary currents $I_{V}(x)$ are known. Formally, it means that the operator $\Lambda^{DtN}$ from Dirichlet data to Neumann data is known:
$\Lambda^{DtN}V(x)=I_{V}(x).$
Problem 1.10.
Calderón problem: It is required to recover a conductivity function $\sigma(x)$ using a known operator $\Lambda^{DtN}.$
Omitting some important details, we briefly describe the idea of the numerical approach to Problem 1.10 following [3].
Algorithm 1.11.
[3]
  • Using the operator $\Lambda^{DtN},$ construct a matrix $M$, which is a response of the general position (well-connected, see [14]) minimal network $\mathcal{E}(G,w)$ on a given graph;
  • Solve Problem 1.8 for $\mathcal{E}$;
  • Using the edge conductivities of $\mathcal{E}$ numerically found out a conductivity $\sigma(x)$.
2. From electrical networks to non-negative Grassmanians
2.1. The Postnikov theory
The central idea of our paper is that the black box problem can be framed as a specific instance of the black box problem within the Postnikov theory, which we briefly describe following [16] and [19].
Definition 2.1.
The totally non-negative Grassmannian $\mathrm{Gr}_{\geq 0}(k,m)$ is a subset of the points in the Grassmannian $\mathrm{Gr}(k,m)$ whose Plücker coordinates $\Delta_{I}$ have the same sign or vanish.
The key result of the Postnikov theory is that each $X\in\mathrm{Gr}_{\geq 0}(k,m)$ can be parametrized by the Lam models.
Definition 2.2.
A planar graph $\Gamma$ together with a weight function $\omega:E(\Gamma)\rightarrow\mathbb{R}_{>0}$ is called a Lam model $N(\Gamma,\omega)$ if the following holds:
  • $\Gamma$ is embedded into a disk $D$, and its nodes are divided on two sets, namely a set of inner nodes $V_{I}$ and a set of boundary nodes $V_{B};$
  • The boundary nodes $V_{B}$ lay on a boundary circle $\partial D$ and are numbered clockwise from $1$ to $n:=|V_{B}|$;
  • The degrees of the boundary nodes are all equal to one;
  • All the vertices of $\Gamma$ are colored by either black or white color and each edge is incident to the vertices of different colors.
Note that when illustrating Lam models, we will omit the colors of boundary nodes and suppress the weights of edges if they are equal to $1.$
Definition 2.3.
A dimer $\Pi$ is a collection of edges of $\Gamma$ such that
  • Each interior vertex is incident to one of the edges in $\Pi$;
  • Boundary vertices may or may not be incident to the edges in $\Pi$.
The weight $\mathrm{wt}(\Pi)$ of a dimer $\Pi$ is a product of all weights of its edges.
Figure 4. A Lam model with $k(\Gamma)=2$ and its dimer $\Pi\in\Pi(15)$ with the weight $\mathrm{wt}(\Pi)$ equals to abab
Consider a Lam model $N(\Gamma,\omega)$ and its arbitrary dimer $\Pi$. It is not difficult to verify that the sum of a number of black boundary nodes are covered by $\Pi$ with a number of white boundary nodes are not covered by $\Pi$ is independent of the choice of dimer $\Pi$. This sum can be calculated as follows:
$k(\Gamma)=\frac{1}{2}\left(n+\sum\limits_{v\in black}(\mathrm{deg}(v)-2)+\sum\limits_{v\in white}(2-\mathrm{deg}(v))\right),$
where summations are over black or white inner nodes of $\Gamma.$ This observation leads to the following definition.
Definition 2.4.
For each $I\subset V_{I}$ such that $|I|=k(\Gamma)$ the boundary measurement is given by the formula:
$\Delta_{I}^{d}=\sum\limits_{\Pi\in\Pi(I)}\mathrm{wt}(\Pi),$
where the summation goes over all dimers $\Pi\in\Pi(I)$ such that the black boundary nodes belonging to $I$ are covered by $\Pi,$ and white nodes belonging to $I$ are not covered by $\Pi$.
Theorem 2.5.
[16] For a Lam model $N(\Gamma,\omega)$ the collection of boundary measurements $\Delta_{I}^{d}$ considered as a set of Plücker coordinates defines a point in $\mathrm{Gr}_{\geq 0}(k(\Gamma),m)$. In particular, the boundary measurements $\Delta_{I}^{d}$ satisfy the Plücker relations.
Moreover, for each point $X\in\mathrm{Gr}_{\geq 0}(k,m)$, there is a Lam model $N(\Gamma,\omega)$ associated with it such that $\Delta_{I}(X)=\Delta^{d}_{I}.$
The second statement of Theorem 2.5 naturally leads to the formulation of the black box problem in the Postnikov theory.
Problem 2.6.
Consider a Lam model $N(\Gamma,\omega)$ on a given graph $\Gamma$ associated with a given point $X\in\mathrm{Gr}_{\geq 0}(k,m)$. It is required to recover a weight function $\omega$ of $N(\Gamma,\omega)$ using Plücker coordinates of $X.$
The possibility of solving Problem 2.6 is revealed by the following result.
Definition 2.7.
Consider a Lam model $N(\Gamma,\omega)$. Then, its oriented median graph $\Gamma_{M}$ is defined as follows:
  • Boundary nodes of $\Gamma_{M}$ coincide with boundary nodes of $N(\Gamma,\omega)$;
  • Inner nodes of $\Gamma_{M}$ are the midpoints of the edges connecting inner nodes of $\Gamma$.
Two nodes of $\Gamma_{M}$ are connected by an edge if the edges of the original graph $\Gamma$ are adjacent. Edges of $\Gamma_{M}$ are oriented as follows:
  • Clockwise around each inner white node of $\Gamma$;
  • Counterclockwise around each inner black node of $\Gamma$.
Since all inner nodes of $\Gamma_{M}$ have degree four, we can correctly define the strands of $\Gamma_{M}$ as the oriented paths which always go straight through any degree four node.
Definition 2.8.
A Lam model $N(\Gamma,\omega)$ is called minimal if it satisfies the following condition:
  • Each strand does not intersect itself;
  • There are no loops apart from loops attached to isolated boundary nodes;
  • Strands do not form oriented lenses, see Fig. 5.
Figure 5. A forbidden oriented lens
Theorem 2.9.
[16],[19] Consider a minimal Lam model $N(\Gamma,\omega)$ on a given graph $\Gamma$ associated with a given point $X\in\mathrm{Gr}_{\geq 0}(k,m)$. Then, a weight function $\omega$ is uniquely recovered ((see Fig. 6)) by the Plücker coordinates of $X,$ up to the gauge transformations.
Figure 6. A gauge transformation in an inner node
Analogous to electrical network theory, the proofs of Theorem 2.9 provide algorithms to solve Problem 2.6. We will mention a few of them:
  • The recursive bridge-lollipop decomposition [16];
  • Solutions to Problem 2.6 for the special Le-diagram models [19], [22];
  • Solution to Problem 2.6 by the generalized chamber ansatz [18].
We shall describe in detail the last approach, since it is very important for our future goal.
Theorem 2.10.
[18],[20] Consider a minimal Lam model $N(\Gamma,\omega)$. Let us label each of its faces $F$ by $I(F)\subset[n]$ according to the following the Scott rule: if a face $F$ lies on the left-hand side of the oriented strand from a source numbered $i$, we have that $i\in I(F)$. This labeling rule ensures that each face is labeled by $k(\Gamma)$ distinct indices.
Definition 2.11.
[18] Consider a matrix $A\in\mathrm{Mat}_{k\times m}(\mathbb{R})$ and denote by $A_{i},\ i\in\{1,\dots,m\}$ its columns. We define the twist $\tau(A)\in\mathrm{Mat}_{k\times m}(\mathbb{R})$ as follows:
[…]
where $\tau(A)_{i},i\in\{1,\dots,m\}$ are the columns of $\tau(A)$; $(\cdot,\cdot)$ is the standard scalar product; and the index operation ++ is taken modulo $m$.
Theorem 2.12.
[18] Consider a minimal Lam model $N(\Gamma,\omega)$ on a given graph $\Gamma$ associated with a point $X\in\mathrm{Gr}_{\geq 0}(k,m)$ defined by a matrix $A.$ Then, a weight function $\omega$ can be found as follows:
  • The first step is to label each face $F$ of $\Gamma$ by $I(F)$ according to the rule from Theorem 2.10;
  • The second step is to find the twist $\tau(A)$ of $A;$
  • Finally we calculate the weight function $\omega$ according to the rule:
    • If an edge $e$ connects two inner nodes, then
      (1) $\omega(e)=\frac{1}{\Delta_{I(F_{1})}\bigr(\tau(A)\bigl)\Delta_{I(F_{2})}\bigr(\tau(A)\bigl)},$
      where the edge $e$ is shared by faces $F_{1}$ and $F_{2}$;
    • If an edge $e$ connects an inner node with a boundary node, then
      (2) $\omega(e)=\frac{1}{\Delta_{I(F)}\bigr(\tau(A)\bigl)},$
      where $F$ is
        The closest counterclockwise which shares the edge $e,$ if $e$ connects a white boundary node;
        The closest clockwise which shares the edge $e,$ if $e$ connects a black boundary node.
Remark 2.13.
For our future goals, we provide the general idea of the proof of Theorem 2.12 following [18]. In fact, each face $F$ of a minimal Lam model $N(\Gamma,\omega)$ uniquely defines a minimal (under the partial order on a set of all dimers with a fixed boundary) dimer $\Pi_{F}$, whose boundary can be calculated according to the rule from Theorem 2.10 i.e. $\Pi_{F}\in\Pi\bigr(I(F)\bigl)$. Using weights $\mathrm{wt}(\Pi_{F})$ of these dimers we can recover the weight function $\omega$ up to the gauge transformations according to the following rule (where $F_{1},F_{2},F$ are defined as in Theorem 2.12):
  • If an edge $e$ connects two inner nodes then
    $\omega(e)=\mathrm{wt}(\Pi_{F_{1}})\mathrm{wt}(\Pi_{F_{2}}),$
  • If an edge $e$ connects an inner node to a boundary node, then
    $\omega(e)=\mathrm{wt}(\Pi_{F}).$
Finally, it remains to prove that the following identity holds:
$\Delta_{I(F)}\bigr(\tau(A)\bigl)\mathrm{wt}(\Pi_{F})=1.$
2.2. The Lam embedding and its parametrization
We now present the cornerstone result that will enable us to apply the Postnikov theory to solve Problem 1.8.
Theorem 2.14.
[6] Consider an electrical network $\mathcal{E}(G,w)\in E_{n}$ with a response matrix $M_{R}(\mathcal{E})=(x_{ij}).$ Then there is an injective map $\mathcal{L}:E_{n}\to\mathrm{Gr}_{\geq 0}(n-1,2n)$ such that $\mathcal{L}(\mathcal{E})$ is a row space of the $(n\times 2n)$ matrix
$\Omega(\mathcal{E})=\left(\begin{array}[]{cccccccc}x_{11}&1&-x_{12}&0&x_{13}&0&\cdots&(-1)^{n}\\ -x_{21}&1&x_{22}&1&-x_{23}&0&\cdots&0\\ x_{31}&0&-x_{32}&1&x_{33}&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots&\vdots\end{array}\right)$
In particular we have:
  • The dimension of the row space of $\Omega(\mathcal{E})$ is equal to n1n-1;
  • Each $n-1\times n-1$ minors of $\Omega(\mathcal{E})$ is non-negative;
  • Plücker coordinates of the point of $\mathcal{L}(\mathcal{E})$ correspond to $n-1\times n-1$ minors of the matrix $\Omega^{\prime}(\mathcal{E})$ obtained from $\Omega(\mathcal{E})$ by deleting the last row.
Remark 2.15.
The non-negativity of the Plücker coordinates of points $\mathcal{L}(\mathcal{E})$ is essentially equivalent to the last property of response matrices described in Theorem 1.3 ((see [5] for more details)).
Due to Theorem 2.5 a point $\mathcal{L}(\mathcal{E})$ associates with a Lam model, which can be canonically constructed with a network $\mathcal{E}$ by the generalized Temperley trick. A model obtained by this trick is essentially constructed by combining a graph $G$ and its dual $G^{*}$. The nodes of both $G$ and $G^{*}$ are colored black. An additional set of nodes introduced at the intersections of edges from $G$ and $G^{*}$ are colored white. Edges originating from $G$ inherit a weight equal to the conductivity of the corresponding edge in $G$, while all new edges are assigned a weight of $1$. More formally:
Definition 2.16.
[15] Given an electrical network $\mathcal{E}(G,w)\in E_{n}$, let us construct its corresponding Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ defined on a graph $\mathcal{E^{T}}$ with weight function $\boldsymbol{\omega}$. The nodes of $\mathcal{E^{T}}$ are defined as follows:
  • If $G$ has boundary nodes $\{1,2,\dots,n\}$, then $\mathcal{E^{T}}$ will have white boundary nodes $\{1,2,\dots,2n\}$, where boundary node $i$ is identified with 2i12i-1 and node $2i$ can be identified with the additional node lied between $i$ and $i+1$;
  • We have a black inner node $b_{v}$ for each inner node $v$ of $G$; a black inner node $b_{F}$ for each interior face $F$ of $G$; a white inner node $w_{e}$ placed at the midpoint of each interior edge $e$ of $G;$
  • For each boundary node $i$ of $G$, we have a black inner node $b_{i}$.
The edges of $N(\mathcal{E^{T}},\boldsymbol{\omega})$ are defined as follows:
  • If $v$ is a node of an edge $e$ in $G$, then $b_{v}$ and $w_{e}$ are joined, and the weight $\boldsymbol{\omega}(e)$ of this edge is equal to the weight $w(e)$ of $e$ in $G$;
  • If $e$ borders $F$, then $w_{e}$ is joined to $b_{F}$ by an edge with weight $1$;
  • The node $b_{i}$ is joined by an edge with weight $1$ to the boundary node 2i12i-1 in $N(\mathcal{E^{T}},\boldsymbol{\omega})$, and $b_{i}$ is also joined by an edge with weight $1$ to $w_{e}$ for any edge $e$ incident to $i$ in $G$;
  • Even boundary nodes $2i$ in $N(\mathcal{E^{T}},\boldsymbol{\omega})$ are joined by an edge with weight $1$ to the face vertex $w_{F}$ of the face $F$ that they lie in.
Figure 7. The triangle-shape network and its Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$
Theorem 2.17.
  • A Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ and $\Omega(\mathcal{E})$ define the same point of $\mathrm{Gr}_{\geq 0}(n-1,2n),$ it was proved in [6];
  • Let us suppose that a network $\mathcal{E}$ is minimal and connected, then a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ is also minimal.
Figure 8. Strands of $G_{M}$ and $\Gamma_{M}$ around a midpoint of an electrical network edge
Proof.
To prove the second statement, it suffices to verify that $N(\mathcal{E^{T}},\boldsymbol{\omega})$ satisfies conditions in Definition 2.8. Indeed, the strands of a median graph $G_{M}$ of a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ are obtained by the parallel transitions (see Fig. 8) of strands of a median graph $\Gamma_{M}$ of a network $\mathcal{E}$. Since $\mathcal{E}$ is minimal, any two strands of $\Gamma_{M}$ intersect at most once and each strand does not intersect itself. Therefore, using the above argument we conclude that:
  • Any two strands of $G_{M}$ can intersect at most once in a midpoint of an inner edge of $N(\mathcal{E^{T}},\boldsymbol{\omega})$;
  • No strand of $G_{M}$ intersect itself.
Thus, any two strands might intersect:
  • Exactly once in a midpoint of an inner edge;
  • Exactly once in a midpoint of an inner edge and exactly once in a boundary node. Since all boundary nodes are white and a midpoint of an inner edge is connected with a white inner node, it is easy to check that this case does not lead to the forbidden lens, see Fig. 5;
  • Exactly once in a midpoint of an inner edge and two times in a boundary nodes. According to the reasoning above, these intersections are not possible, since they contradict to the orientation rule of $G_{M};$
  • Two times in a boundary nodes. This case also does not lead to the forbidden lens.
Therefore, any two strands do not form a forbidden lens, it ends the proof. ∎
3. The solution to the black problem
3.1. Conductivity recovering via the generalized chamber ansatz
The generalized Temperley trick and Theorems 2.9 and 2.17 allow us to reduce Problem 1.8 to a particular case of Problem 2.6:
Proposition 3.1.
Consider a minimal electrical network $\mathcal{E}(G,w)$. Then, the recovery of a conductivity function $w$ by a response matrix $M_{R}(\mathcal{E})$ is equivalent to the recovery of a weight function $\omega$ of a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ by a matrix $\Omega(\mathcal{E})$.
As a consequence of Theorems 2.9 and 2.12, we obtain that, for any minimal network, we can reconstruct a weight function $\boldsymbol{\omega}^{\prime}$ for a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$. This function is equal to the original weight function $\boldsymbol{\omega}$ from Definition 2.16 up to gauge transformations. However, explicitly determining the sequence of gauge transformations necessary to transform $\boldsymbol{\omega}^{\prime}$ into $\boldsymbol{\omega}$ is often difficult. Fortunately, there are specific coordinates on $N(\mathcal{E^{T}},\boldsymbol{\omega})$ that remain invariant under such gauge transformations:
Lemma 3.2.
[19] We define the weight of each face $F$ of $N(\mathcal{E^{T}},\boldsymbol{\omega})$ as follows:
$O(F)=\frac{\prod\limits_{e\in F,\ e=wb}\boldsymbol{\omega}(e)}{\prod\limits_{e\in F,\ e=bw}\boldsymbol{\omega}(e)},$
where $F$ is considered to be clockwise edge oriented; the numerator is the product taken over all from white to black edges of $F$ and the denominator is the product taken over all from black to white edges of $F$.
The face weight $O(F)$ does not depend on the gauge transformations.
Figure 9. A bridge and a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ around it
Figure 10. A spike and a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ around it
Figure 11. Two adjacent edges and a Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ around them
Using the coordinates introduced in Lemma 3.2, we can formulate a complete analog of the chamber ansatz (see [1], [18]) for the conductivities of electrical networks:
Lemma 3.3.
Consider a minimal electrical network $\mathcal{E}(G,w)$ and its edge $e$. Then, the following holds:
  • If an edge $e$ is a bridge ((i.e. an edge which connects a boundary node numbered by $i$ with a boundary node numbered by $i+1)$ its conductivity can be recovered by the formula
    (3) $w(e)=\dfrac{\Delta_{1}\Delta_{3}}{\Delta_{2}\Delta_{4}}$
    or, alternatively, by the formula
    (4) $w(e)=\dfrac{\Delta_{1}}{\Delta_{5}};$
  • If an edge $e$ is a spike ((i.e. an edge which connects a boundary node with an unique inner node)) can be recovered by the formula
    (5) $w(e)=\dfrac{\Delta_{2}\Delta_{4}}{\Delta_{1}\Delta_{3}}$
    or, alternatively, by the formula
    (6) $w(e)=\dfrac{\Delta_{5}}{\Delta_{1}};$
  • Let us consider a face $F$ and its two adjacent edges $e_{1}$ and $e_{2}$ ((see Fig. 11)), then the ratio of their conductivities is equal to
    (7) $\dfrac{w(e_{1})}{w(e_{2})}=\dfrac{\Delta_{2}\Delta_{4}}{\Delta_{1}\Delta_{3}},$
where $F_{i}$ are the faces labeled by $\Delta_{i}$ as it is shown in Fig. 9, Fig. 10 and Fig. 11; and by $\Delta_{i}$ we denote
$\Delta_{i}:=\Delta_{I(F_{i})}\Bigr(\tau\bigr(\Omega^{\prime}(\mathcal{E})\bigl)\Bigl).$
Proof.
Let us prove Formula (3). Indeed, let us consider a face $F_{1}$, then
$O(F_{1})=\dfrac{\boldsymbol{\omega}(e_{1})\boldsymbol{\omega}(e_{3})}{\boldsymbol{\omega}(e_{2})\boldsymbol{\omega}(e_{4})}=a.$
On the other hand, using Theorem 2.12 we have that
$O(F_{1})=\dfrac{\boldsymbol{\omega}^{\prime}(e_{1})\boldsymbol{\omega}^{\prime}(e_{3})}{\boldsymbol{\omega}^{\prime}(e_{2})\boldsymbol{\omega}^{\prime}(e_{4})}=\dfrac{\Delta_{1}\Delta_{3}}{\Delta_{2}\Delta_{4}},$
where up to gauge transformations $\boldsymbol{\omega}^{\prime}=\boldsymbol{\omega}$ and
$\boldsymbol{\omega}^{\prime}(e_{1})=\dfrac{1}{\Delta_{4}},\ \boldsymbol{\omega}^{\prime}(e_{2})=\dfrac{1}{\Delta_{1}\Delta_{4}},\ \boldsymbol{\omega}^{\prime}(e_{3})=\dfrac{1}{\Delta_{4}\Delta_{2}},\ \boldsymbol{\omega}^{\prime}(e_{4})=\dfrac{1}{\Delta_{3}}.$
Since the value of $O(F_{1})$ is independent on the gauge transformations, we obtain Formula (3). The remaining formulas can be derived in a similar way, which completes the proof. ∎
Based on Lemma 3.3, we propose the following algorithm to solve Problem 1.8.
Algorithm 3.4.
  • Given a response matrix $M_{R}(\mathcal{E})$ of a minimal network $\mathcal{E}$ construct the matrix $\Omega^{\prime}(\mathcal{E})$;
  • Calculate $\tau\bigl(\Omega^{\prime}(\mathcal{E})\bigr);$
  • Label each face $F$ of $N(\mathcal{E^{T}},\boldsymbol{\omega})$ by $n-1\times n-1$ minors $\Delta_{F(I)}$ of a $\tau\bigl(\Omega^{\prime}(\mathcal{E})\bigr)$;
  • Using Formulas (3)–(4), (5)–(6) and (7) recover a conductivity function $w.$
Proof.
The correctness of the proposed algorithm is derived from the following statements:
  • Each minimal electrical network has at least one bridge or spike (see [8]), therefore we can extend bridges and spikes conductivities to conductivities of remaining edges by Formula (7);
  • The uniqueness of this extension defined by Formula (7) follows from Theorem 2.9 and the second part of Theorem 2.17.
Figure 12. The star-shape network $\mathcal{E}$
Remark 3.5.
As demonstrated by Equations (3)–(7), the conductivity function $w$ can be recovered in multiple ways. Furthermore, every minimal electrical network contains at least three bridges and spikes in total, see [8]. Consequently, the process of recovering the conductivity function $w$ leads to a multitude of non-trivial identities, which of course follow from the special form of the weight function $\boldsymbol{\omega}$ of $N(\mathcal{E^{T}},\boldsymbol{\omega})$. These identities may hint at the existence of integrable structures (see [2]) related to electrical networks.
Let us consider a couple of examples.
Example 3.6.
Consider a network $\mathcal{E}$ as it is shown in Fig. 12. Its response matrix has the form:
$M_{R}(\mathcal{E})=\left(\begin{array}[]{cccccccc}\frac{ca+cb}{a+b+c}&-\frac{ca}{a+b+c}&-\frac{cb}{a+b+c}\\ -\frac{ca}{a+b+c}&\frac{ca+ab}{a+b+c}&-\frac{ab}{a+b+c}\\ -\frac{cb}{a+b+c}&-\frac{ab}{a+b+c}&-\frac{cb+ab}{a+b+c}\end{array}\right).$
Then, $\Omega_{3}(\mathcal{E})$ and $\Omega^{\prime}_{3}(\mathcal{E})$ equal to:
$\Omega_{3}(\mathcal{E})=\left(\begin{array}[]{cccccccc}\frac{ca+cb}{a+b+c}&1&\frac{ca}{a+b+c}&0&\frac{-cb}{a+b+c}&-1\\ \frac{ca}{a+b+c}&1&\frac{ca+ab}{a+b+c}&1&\frac{ab}{a+b+c}&0\\ \frac{-cb}{a+b+c}&0&\frac{ab}{a+b+c}&1&\frac{cb+ab}{a+b+c}&1\end{array}\right),\ \Omega^{\prime}_{3}(\mathcal{E})=\left(\begin{array}[]{cccccccc}\frac{ca+cb}{a+b+c}&1&\frac{ca}{a+b+c}&0&\frac{-cb}{a+b+c}&-1\\ \frac{ca}{a+b+c}&1&\frac{ca+ab}{a+b+c}&1&\frac{ab}{a+b+c}&0\par\end{array}\right).$
By direct computation we can verify that all $2\times 2$ minors of $\Omega^{\prime}_{3}(\mathcal{E})$ do not vanish, therefore
$\tau\bigl(\Omega^{\prime}_{3}(\mathcal{E})\bigr)=\left(\begin{array}[]{cccccccc}\frac{a+b+c}{bc}&\frac{c+b}{b}&\frac{a+b+c}{ac}&\frac{a}{c}&0&-1\\ -\frac{a+b+c}{bc}&-\frac{c}{b}&0&1&\frac{a+b+c}{ab}&\frac{a+b}{a}\end{array}\right).$
For instance, $\tau(A)_{6}$ is defined by the following conditions: $(\tau(A)_{6},A_{6})=1$ and $(\tau(A)_{6},A_{6+1\ \text{mod}\ 6})=(\tau(A)_{6},A_{1})=0$ (since $A_{6}\notin\text{span}(A_{1})$).
Let us label each face of $N(\mathcal{E^{T}},\boldsymbol{\omega})$ according to the rule defined in Theorem 2.10, as it is shown in Fig. 13. Since the edge $e_{1}$ is a spike, we obtain the following identities:
(8) $w(e_{1})=\dfrac{\Delta_{26}}{\Delta_{16}}=\dfrac{\frac{a+b+c}{a}}{\frac{a+b+c}{ca}}=c,$
or alternatively
(9) $w(e_{1})=\dfrac{\Delta_{56}\Delta_{46}}{\Delta_{16}\Delta_{45}}=\dfrac{\frac{a+b+c}{ab}\frac{a+b+c}{c}}{\frac{a+b+c}{ca}\frac{a+b+c}{cb}}=c.$
Figure 13. The Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega})$ and the face labeling
Figure 14. $G_{5}$ and its associated Lam model. The boundary circles are not drawn
Example 3.7.
Let us consider a series of the general position (well-connected, see [14]) minimal electrical networks $\mathcal{E}(G_{4m+1},w)$ defined as follows:
  • The boundary nodes lie on a boundary circle of radius $m+1$, centered at $(0,0)$;
  • The nodes of $G_{4m+1}$ are the points $(i,j)$ for integers $i$ and $j$ with $0<i\leq m+1$ and $1\leq j\leq 4m+1$. The radial edges are the radial line segments joining $(i,j)$ to $(i+1,j)$ for each $0<i\leq m$ and $1\leq j\leq 4m+1$. The circular edges are the circular arcs joining $(i,j)$ to $(i,j+1)$ for each $1\leq i\leq m$ and each $1\leq j\leq 4m+1$. The graph $G_{4m+1}$ has $2m(4m+1)$ edges and (m1)(4m+1)(m-1)(4m+1) nodes;
  • The boundary nodes of $G_{4m+1}$ are the points $v_{j}=(m+1,j)$, for $j=1,\ldots,4m+1$, with the convention that $v_{0}=v_{4m+1}$.
The significance of these electrical networks follows from their role as duals to the networks utilized in numerical solutions to the Calderón problem, see [3] and also [11, Theorem $3.7$].
Based on the specific shape of $\mathcal{E}(G_{4m+1},w)$, we can propose the following way to solve Problem 1.8 to them:
  • Label each face of $N(\mathcal{E^{T}},\boldsymbol{\omega})$ according to the rule described in [5];
  • Recover spikes conductivities;
  • Using a spike conductivity and Formula (7), extend conductivities of the remaining edges in each part, see Fig. 15 and 16:
    $a_{1}=\dfrac{\Delta_{1}\Delta_{2}}{\Delta^{\prime}_{1}\widetilde{\Delta}_{1}},\ \dfrac{a_{i}}{b_{i}}=\dfrac{\Delta_{2i-1}\Delta_{2i+1}}{\Delta^{\prime}_{2i}\widetilde{\Delta}_{2i}},\ \dfrac{b_{i}}{a_{i+1}}=\dfrac{\Delta^{\prime}_{2i+1}\widetilde{\Delta}_{2i+1}}{\Delta_{2i}\Delta_{2i+2}}.$
Figure 15. The part
Figure 16. A Lam model across the part
3.2. Conductivity recovering via groves
Now, we reveal some combinatorial aspects of Algorithm 3.4. Particularly, we demonstrate that conductivities can be recovered by weights of the special groves.
Definition 3.8.
A grove $\mathcal{G}$ on a graph $G$ of an electrical network $\mathcal{E}(G,w)$ is a spanning forest, i.e. an acyclic subgraph of $G$ covering all nodes of $\mathcal{E}$, such that each connected component of $\mathcal{G}$ contains at least one boundary node.
A weight $wt(\mathcal{G})$ of the grove $\mathcal{G}$ is equal to $wt(\mathcal{G})=\prod\limits_{e\in\mathcal{G}}w(e).$
Theorem 3.9.
[15] Consider an electrical network $\mathcal{E}(G,w)$ and its associated Lam model $N(\mathcal{E^{T}},\boldsymbol{\omega}).$ Then, for each $I\subset\{1,\dots,n\},\ |I|=n-1$, there is the natural weight preserving bijection $\mathcal{F}_{I}$ between the subset of dimers $\Pi\in\Pi(I)$ on $N(\mathcal{E^{T}},\boldsymbol{\omega})$ ((i.e. a subset of dimers, which do not cover boundary nodes belonging to $I$, see Definition 2.4)) and its corresponding subset of groves on $\mathcal{E}(G,w)$ such that $wt(\Pi)=wt\bigl(\mathcal{F}_{I}(\Pi)\bigr).$
Due to Theorem 3.9, the Plücker coordinates of $\mathcal{L}(\mathcal{E})$ can be rewritten in terms of grove weights.
Theorem 3.10.
[6] Consider an electrical network $\mathcal{E}(G,w)\in E_{n}$. Then, the following identity holds:
$\Delta_{I}\bigl(\Omega(\mathcal{E})\bigr)=\dfrac{\sum\limits_{\Pi\in\Pi(I)}wt(\Pi)}{L_{unc}}=\dfrac{\sum\limits_{\Pi\in\Pi(I)}wt\bigl(\mathcal{F}_{I}(\Pi)\bigr)}{L_{unc}}=\dfrac{\sum\limits_{\mathcal{G}|I}wt\bigl(\mathcal{G}\bigr)}{L_{unc}},$
where $L_{unc}$ is the partition function of the weights of all groves with exactly $n$ connected components; and the last sum goes over all groves, which are concordant with II, see [15] for more details.
Theorem 3.9 and Remark 2.13 also provide additional interpretations of the formulas presented in Lemma 3.3.
Proposition 3.11.
Consider a minimal electrical network $\mathcal{E}(G,w)$. Then, the following Plücker coordinate equality holds:
$\Delta_{I(F)}\Bigr(\tau\bigr(\Omega^{\prime}(\mathcal{E})\bigl)\Bigl)=\dfrac{1}{wt(\mathcal{G}_{F})},$
where $\mathcal{G}_{F}:=\mathcal{F}_{I}(\Pi_{F})$ is the minimal grove corresponded to the minimal dimer $\Pi_{F},$ see Remark 2.13.
Remark 3.12.
Moreover, if a face $F$ of $N(\mathcal{E^{T}},\boldsymbol{\omega})$ is a boundary face (i.e. a face which is bounded by the boundary circle of $N(\mathcal{E^{T}},\boldsymbol{\omega})$), then $\Pi_{F}$ is a unique dimer such that $\Pi\in I_{F}(\Pi)$, see [18]. Therefore, we obtain the following Plücker coordinate equality for the boundary face:
(10) $\Delta_{I(F)}\bigr(\Omega^{\prime}(\mathcal{E})\bigl)=wt(\mathcal{G}_{F}).$
Proposition 3.13.
Combining Proposition 3.11 with Lemma 3.3, we conclude that the conductivity function of each minimal electrical network can be recovered by weights of the minimal groves.
Let us provide an example:
Figure 17. Minimal groves corresponding to faces labeled by $26$ and $16$
Example 3.14.
Consider a network $\mathcal{E}$ as it is shown in Fig. 12. Then, Formula (8) can be rewritten as follows:
$w(e_{1})=\dfrac{\frac{1}{wt(\mathcal{G}_{26})}}{\frac{1}{wt(\mathcal{G}_{16})}}=\dfrac{\frac{1}{a}}{\frac{1}{ac}}=c.$
Proposition 3.13 naturally leads to the following question:
Problem 3.15.
In [18] minimal dimers $\Pi_{F}$ of a minimal Lam model $N(\Gamma,\omega)$ were defined using strands of an oriented median graph $\Gamma_{M}$ of $N(\Gamma,\omega).$ Is there a similar approach for constructing minimal groves $\mathcal{G}_{F}$ of a minimal electrical network $\mathcal{E}(G,w)$ with strands of median graph $G_{M}$ of $\mathcal{E}(G,w)$?
3.3. Conductivity recovering via effective resistances
In this section, we present a solution to an alternative formulation of Problem 1.8 that, beyond the motivations outlined in Section 1.3, also finds application in the theory of phylogenetic networks, as demonstrated in [9], [10] and [11].
Definition 3.16.
Let $\mathcal{E}(G,w)\in E_{n}$ be a connected electrical network, and let the boundary voltages $U=(U_{1},\dots,U_{n})$ be such that
$M_{R}(\mathcal{E})U=-e_{i}+e_{j},$
where $e_{k},\ k\in\{1,\dots,n\}$ is the standard basis of $\mathbb{R}^{n}.$
Let us define the effective resistance $R_{ij}$ between nodes $i$ and $j$ as follows $|U_{i}-U_{j}|:=R_{ij}.$ Effective resistances are well-defined and $R_{ij}=R_{ji}$.
We will organize the effective resistances $R_{ij}$ in an effective resistance matrix $R_{\mathcal{E}}$ by setting $R_{ii}=0$ for all $i$.
Problem 3.17.
Consider an electrical network $\mathcal{E}(G,w)$ on a given graph $G.$ The conductivity function $w$ is required to be recovered by a known effective resistance matrix $R_{\mathcal{E}}$.
The following theorem reveals the possibility of solving the described version of the black box problem:
Theorem 3.18.
The conductivity function $w$ of a minimal electrical network $\mathcal{E}(G,w)\in E_{n}$ can be uniquely recovered by a known effective resistance matrix $R_{\mathcal{E}}$.
We provide a constructive proof of Theorem 3.18, which will be absolutely similar to the proof of Theorem 1.9 due to the following result:
Theorem 3.19.
[4] Let $\mathcal{E}(G,w)\in E_{n}$ be a connected network. Then, using its effective resistance matrix $R_{\mathcal{E}}$, we define a point in Gr(n1,2n)Gr(n-1,2n) associated with it as the row space of the matrix:
$\Omega_{R}(\mathcal{E})=\left(\begin{matrix}1&m_{11}&1&-m_{12}&0&m_{13}&0&\ldots\\ 0&-m_{21}&1&m_{22}&1&-m_{23}&0&\ldots\\ 0&m_{31}&0&-m_{32}&1&m_{33}&1&\ldots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{matrix}\right),$
where
$m_{ij}=-\frac{1}{2}(R_{i,j}+R_{i+1,j+1}-R_{i,j+1}-R_{i+1,j}).$
Then, $\Omega_{R}(\mathcal{E})$ defines the same point of $Gr_{\geq 0}(n-1,2n)$ as the point $\mathcal{L}(\mathcal{E}),$ see Theorem 2.14.
Example 3.20.
Consider a network as it is shown in Fig. 18. Then, its effective resistance matrix has the form
$R_{\mathcal{E}}=\begin{pmatrix}0&3&3&2\\ 3&0&2&3\\ 3&2&0&3\\ 2&3&3&0\end{pmatrix}$,
and the matrix $\Omega^{\prime}_{R}(\mathcal{E})$ obtained from $\Omega_{R}(\mathcal{E})$ by deleting its last row has the form
$\Omega^{\prime}_{R}(\mathcal{E})=\begin{pmatrix}1&3&1&1&0&-1&0&1\\ 0&1&1&2&1&1&0&0\\ 0&-1&0&1&1&3&1&1\end{pmatrix}$.
Figure 18. All conductivities are equal to 11
Since the matrices $\Omega_{R}^{\prime}(\mathcal{E})$ and $\Omega^{\prime}(\mathcal{E})$ define the same point in $Gr_{\geq 0}(n-1,2n)$, we immediately derive the following algorithm for solving Problem 3.17:
Algorithm 3.21.
  • By a given effective resistance matrix $R_{\mathcal{E}}$ construct a matrix $\Omega_{R}^{\prime}(\mathcal{E})$;
  • Calculate $\tau\bigl(\Omega_{R}^{\prime}(\mathcal{E})\bigr);$
  • Do the last two steps of Algorithm 3.4.
4. The black box problem as the Berenstein-Fomin-Zelevinsky problem
In this section, we discuss the connection between Problem 1.8 and the classical theory of totally non-negative matrices. To achieve this goal, we primarily describe an embedding of electrical networks into the positive part of the Lagrangian Grassmannian.
Lemma 4.1.
[6] Let us define the subspace
$V=\{v\in\mathbb{R}^{2n}|\sum_{i=1}^{n}(-1)^{i}v_{2i}=0\,\sum_{i=1}^{n}(-1)^{i}v_{2i-1}=0\}.$
Then, the rows of the matrix $\Omega(\mathcal{E})$ belong to the subspace $V$. Thus the matrix $\Omega(\mathcal{E})$ defines a point of Gr(n1,V).Gr(n-1,V).
By choosing a particular basis for the subspace $V$, we obtain an embedding of electrical networks into a Lagrangian Grassmannian LG(n1,V)LG(n-1,V) with respect to a symplectic form that depends explicitly on the chosen basis of $V$. Moreover, specific choices (see [6], [21]) of basis $V$ allow us to construct embeddings of electrical networks into the positive part of the Lagrangian Grassmannians $LG_{\geq 0}(n-1,V):=LG(n-1,V)\cap Gr_{\geq 0}(n-1,V)$:
Theorem 4.2.
[12],[21] When $n$ is odd, there exists a special basis of the subspace $V$ such that expanding the rows of the matrix $\Omega(\mathcal{E})$ with respect to this basis gives us a matrix $\overline{\Omega}(\mathcal{E})$ that defines a point of $LG_{\geq 0}(n-1,V)$ with respect to the non-degenerate symplectic form:
$\Lambda^{\prime}_{2n-2}=\left(\begin{array}[]{cc}\Lambda_{n}&0\\ 0&-\Lambda_{n}\\ \end{array}\right),$
where $\Lambda_{n}$ is defined as in (11).
An electrical network is called well-connected if it is equivalent to the standard network, see [14] and Fig. 19.
Theorem 4.3.
[12],[21] Consider well-connected electrical networks with an odd number of boundary nodes. Then, the matrix $\overline{\Omega}(\mathcal{E})$ and a matrix of the form described below represent the same point in $LG_{\geq 0}(n-1,V)$:
$\bigl(\overline{\mathrm{Id}},A(\mathcal{E})\bigr)=\left(\begin{array}[]{ccccccccc}0&\dots&0&0&1&a_{11}&a_{12}&a_{13}&\dots\\ 0&\dots&0&-1&0&a_{21}&a_{22}&a_{23}&\dots\\ 0&\dots&1&0&0&a_{31}&a_{32}&a_{33}&\dots\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\end{array}\right),\ \overline{Id}\in\mathrm{Mat}_{n-1\times n-1}(\mathbb{R}).$
Definition 4.4.
An invertible matrix $A\in\mathrm{Mat}_{n\times n}(\mathbb{R})$ is called totally non-negative (positive) if all its minors are non-negative (positive).
Theorem 4.5.
[12],[21] Consider a well-connected network $\mathcal{E}$ with an odd number of boundary nodes and a standard network $\mathcal{E^{\prime}}$ that is equivalent to $\mathcal{E}$. Then, the matrix $A(\mathcal{E})=(a_{ij})\in\mathrm{Mat}_{n-1\times n-1}(\mathbb{R})$ defined below is a totally positive symplectic matrix with respect to the skew-symmetric non-degenerate form:
(11) $\Lambda_{n}=\left(\begin{array}[]{ccccccccc}0&1&0&0&\dots\\ -1&0&-1&0&\dots\\ 0&1&0&1&\dots&\\ \vdots&\vdots&\vdots&\ddots&\vdots\end{array}\right),\ A(\mathcal{E})\Lambda_{n}A(\mathcal{E})^{t}=\Lambda_{n}.$
Moreover, the matrix $A(\mathcal{E})$ can be decomposed into the product of the elementary totally non-negative symplectic generators $u_{i,i+1}(t)=E+t(E_{i+1,i}+E_{i-1,i})$:
(12) $A(\mathcal{E})=\prod\limits_{1\leq i<j\leq n}u_{j-i,j-i+1}(t_{ij}),$
where $E$ is the identity matrix, $E_{ij}$ is the matrix unity and $E_{n,n-1}=E_{0,1}=0$; the product is taken over all pairs $(i,j)$ ordered in reverse lexicographical order; and
$t_{ij}=w_{ij}^{(-1)^{i+j}},$
here $w_{ij}$ is a conductivity of an edge of the network $\mathcal{E^{\prime}}$ lying in intersection strands numbered $i$ and $j$, see Fig. 20.
Remark 4.6.
The decomposition (12) is compatible with the bridge-spike decomposition, see [6, Section $8.1$]. We suppose that based on this observation it is possible to extend Theorem 4.2 and Theorem 4.5 to any network.
Figure 19. Standard networks, boundary nodes are bold
Figure 20. The edge enumeration
Theorem 4.5 provides an explicit link between electrical network theory and the theory of totally non-negative matrices. The cornerstone result of the last one is that:
Theorem 4.7.
[26] Each totally non-negative matrix can be decomposed into a product of elementary generators, i.e. Jacobi matrices $x_{i}(t)=E+tE_{i,i+1},$ $y_{i}(t)=E+tE_{i+1,i}$ and a diagonal matrix $d=\mathrm{diag}(d_{1},\dots,d_{n})$ with all non-negative parameters $t,d_{i}$.
Example 4.8.
The following holds:
$x_{1}(t)=\left(\begin{array}[]{ccc}1&t&0\\ 0&1&0\\ 0&0&1\end{array}\right),\ y_{2}(t)=\left(\begin{array}[]{ccc}1&0&0\\ 0&1&0\\ 0&t&1\end{array}\right).$
The totally non-negative symplectic generators $u_{i,i+1}(t)\in\mathrm{Mat}_{n-1\times n-1}(\mathbb{R})$ can be decomposed as follows:
$u_{1,2}(t)=y_{1}(t),\ u_{i,i+1}(t)=y_{i+1}(t)x_{i-1}(t),\ u_{n,n+1}(t)=x_{n-2}(t).$
Figure 21. The elementary generators and their corresponded elementary Lam models
Theorem 4.7 gives rise to the following natural problem, which, building upon the works [1], [17], we will call the Berenstein-Fomin-Zelevinsky problem:
Problem 4.9.
Consider a totally non-negative matrix $A$ and its given decomposition in a product of elementary generators. It is required to recover the parameters of elementary generators.
According to Theorem 4.5, Problem 1.8 for well-connected networks with an odd number of boundary nodes can be considered as a special case of Problem 4.9, which, in turn, is closely related to Problem 2.6. This relationship is established by the following trick:
Theorem 4.10.
[19] Let $A\in\mathrm{Mat}_{n\times n}(\mathbb{R})$ be a totally non-negative matrix with a known decomposition into a product of the elementary generators. Consider a point $X(A)\in Gr_{\geq 0}(n,2n)$ which has the following form:
$X(A)=(\overline{\mathrm{Id}},A)=\left(\begin{array}[]{ccccccccc}0&\dots&0&0&1&a_{11}&a_{12}&a_{13}&\dots\\ 0&\dots&0&-1&0&a_{21}&a_{22}&a_{23}&\dots\\ 0&\dots&1&0&0&a_{31}&a_{32}&a_{33}&\dots\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots&\vdots&\ddots&\vdots\end{array}\right),$
The point $X(A)$ is defined by a Lam model that is compatible with the decomposition of $A$ and can be built by concatenating elementary Lam models. Each elementary Lam model corresponds to an elementary generator, as it is shown in Fig. 21 ((here, we use the non-standard notation for the boundary nodes numbering, which relates to an ordinary with the substitution $n\to 1,\dots,1\to n,1^{\prime}\to n+1,\dots,n^{\prime}\to 2n)$.
Example 4.11.
Consider the standard electrical network with $3$ boundary nodes. Using Theorem 4.2, we obtain that
$A(\mathcal{E})=u_{12}(t_{23})u_{23}(t_{13})u_{12}(t_{12})$
and the point $X\bigl(A(\mathcal{E})\bigr)$ is defined by the Lam model, as it is shown in Fig. 22.
Figure 22. The Lam model associates with the standard electrical network with $3$ boundary nodes
Note that the Lam models associated with points $X\bigl(A(\mathcal{E})\bigr)$ are generally not minimal. For instance, even the Lam model depicted in Fig. 22 is not minimal. Consequently, the recovery technique outlined in Section 2.1 cannot be directly applied to solve the Berenstein-Fomin-Zelevinsky problem for totally non-negative symplectic matrices $A(\mathcal{E})$. However, we believe that articulating the connection described above between the black box problem and the Berenstein-Fomin-Zelevinsky problem might be useful. In particular, we expect the following to occur:
Conjecture 4.12.
The following algorithm for solving Problem 1.8 is correct:
  • Using a response matrix $M_{R}(\mathcal{E})$ of a well-connected network $\mathcal{E}$ with an odd number of boundary node, construct the matrix $A(\mathcal{E})$;
  • Using the decomposition (12) and Theorem 4.10, construct a Lam model. Using the Postnikov transformations (see [16]), into a minimal Lam model (we expect that is multiple simple, canonical ways exist for this reduction);
  • Using Theorem 2.12, reconstruct the weights of the minimal Lam model. Reconstruct the weights of the original Lam model from these;
  • Using the found weights and the physical sense of parameters of the decomposition (12), construct a standard network $\mathcal{E^{\prime}}$ equivalent to $\mathcal{E}$. Transform $\mathcal{E^{\prime}}$ into $\mathcal{E}$ using star-triangle transformations.